A sample of vinegar was titrated with 1.054M NaOH. 21.6mL of NaOH were required to reach the endpoint. Find the mass of HC2H3O2 in grams that were in the sample of vinegar.

Hint: answer included 3 sig figs
Answer #1

Balanced chemical equation is:

NaOH + C2H4O2 ---> NaC2H3O2 + H2O

lets calculate the mol of NaOH

volume , V = 21.6 mL

= 2.16*10^-2 L

use:

number of mol,

n = Molarity * Volume

= 1.054*2.16*10^-2

= 2.277*10^-2 mol

According to balanced equation

mol of C2H4O2 reacted = (1/1)* moles of NaOH

= (1/1)*2.277*10^-2

= 2.277*10^-2 mol

This is number of moles of C2H4O2

Molar mass of C2H4O2,

MM = 2*MM(C) + 4*MM(H) + 2*MM(O)

= 2*12.01 + 4*1.008 + 2*16.0

= 60.052 g/mol

use:

mass of C2H4O2,

m = number of mol * molar mass

= 2.277*10^-2 mol * 60.05 g/mol

= 1.367 g

Answer: 1.37 g